Calculate the ΔH of the following problems:





1) How many joules are required to raise the temperature of 450 mL of water from 30°C to 100°C?





2) The specific heat of copper metal is 0.385 J/g °K. How many kJ of heat are needed to raise the temperature of 1.1 kg block of copper from 27.5°C to 50°C?

1
Expert's answer
2022-05-09T17:03:05-0400

1) Q = 450 • 4.18 • 70 = 131.6 kJ

2)M=1.1kg,C=0.385J/g °K, change in temp=50-27.5=22.5°C=295.5 K

Q=1.1x0.385x295.5K = 0.125 kJ